# Meta-Learning Driven Movable-Antenna-assisted Full-Duplex RSMA for Multi-User Communication: Performance and Optimization

### v2

@@ 11             $\textbf{u}^{(0,j)}=\textbf{u}^{(0,1)}$;
14                   $R_{\textbf{P}}^{(i-1,j)}=\max\sum_{d\in\mathcal{D}}R_{d}(\hat{\textbf{W}},\hat{\textbf{u}},\hat{\textbf{c}})+\sum_{u\in\mathcal{U}}R_{u}(\textbf{P}^{(i-1,j)},\hat{\textbf{u}},\hat{\textbf{Z}})$;
@@ 19             Apply projection if applicable as (41)-(44)
21                   $R_{\textbf{W}}^{(i-1,j)}=\max\sum_{d\in\mathcal{D}}R_{d}({\textbf{W}^{(i-1,j)}},\hat{\textbf{u}},\hat{\textbf{c}})+\sum_{u\in\mathcal{U}}R_{u}\hat{(\textbf{P}},\hat{\textbf{u}},\hat{\textbf{Z}})$;
@@ 26             Apply projection as (47)
29                   $R_{\textbf{Z}}^{(i-1,j)}=\max\sum_{d\in\mathcal{D}}R_{d}({\hat{\textbf{W}}},\hat{\textbf{u}},\hat{\textbf{c}})+\sum_{u\in\mathcal{U}}R_{u}\hat{(\textbf{P}},\hat{\textbf{u}},{\textbf{Z}^{(i-1,j)}})$;
@@ 33            $\textbf{Z}^{*}=\textbf{Z}^{(N_{i},j)}$;
36                   $R_{\textbf{C}}^{(i-1,j)}=\max\sum_{d\in\mathcal{D}}R_{d}({\hat{\textbf{W}}},\hat{\textbf{u}},\textbf{c}^{(i-1,j)})+\sum_{u\in\mathcal{U}}R_{u}\hat{(\textbf{P}},\hat{\textbf{u}},\hat{\textbf{Z}}),$;
@@ 40            $\textbf{C}^{*}=\textbf{C}^{(N_{i},j)}$;
43                   $R_{\textbf{u}}^{(i-1,j)}=\max\sum_{d\in\mathcal{D}}R_{d}({\hat{\textbf{W}}},\textbf{u}^{(i-1,j)},\hat{\textbf{c}})+\sum_{u\in\mathcal{U}}R_{u}\hat{(\textbf{P}},\textbf{u}^{(i-1,j)},\hat{\textbf{Z}})$;
@@ 48             Calculate $\mathcal{L}^{j}$ as (54);
49             ¯=¯$\bar{\mathcal{L}}=\bar{\mathcal{L}}+\mathcal{L}^{j}$;
50             If $-\mathcal{L}^{j}>MAX$
51              MAX=$-\mathcal{L}^{j}$;
52            $\textbf{P}_{opt}=\textbf{P}^{*}$;
53            $\textbf{W}_{opt}=\textbf{W}^{*}$;
54            $\textbf{Z}_{opt}=\textbf{P}^{*}$;
55            $\textbf{C}_{opt}=\textbf{C}^{*}$;
56            $\textbf{u}_{opt}=\textbf{u}^{*}$;
@@
59      Calculate ¯$\bar{\mathcal{L}}=\frac{1}{N_{o}}\sum_{j=1}^{N_{o}}\mathcal{L}^{j}$;
@@ V Simulation results
V-A Simulation Settings
In the simulation, we assume two DL UEs and two UL UEs. The UEs are randomly and uniformly distributed in a cell centered on the FD BS located at (0,0) in an area of $200m\times 200m$. The FD BS is equipped with an equal number of transmit and receive antennas, i.e., $N_{T}=N_{R}$. Following [24], we adopt the same channel where we assume that $L_{t}^{BS}=L_{r}^{BS}=L_{t}^{DL}=L_{t}^{u}=L_{t}^{d}=L_{r}^{SI}=L_{t}^{SI}\equiv L$. In this way, every PRM $\boldsymbol{\Sigma}_{d}$, $\boldsymbol{\Sigma}_{u}$, $\boldsymbol{\Sigma}_{SI}$, $\boldsymbol{\Sigma}_{u_{d}}$ becomes a diagonal matrix where $\boldsymbol{\Sigma_{d}}=\textrm{diag}\{\sigma_{1,1},\sigma_{1,1},\dots,\sigma_{L,L}\}$, $\boldsymbol{\Sigma_{u}}=\textrm{diag}\{\sigma_{1,1},\sigma_{1,1},\dots,\sigma_{L,L}\}$, $\boldsymbol{\Sigma_{SI}}=\textrm{diag}\{\sigma_{1,1},\sigma_{1,1},\dots,\sigma_{L,L}\}$, $\boldsymbol{\Sigma_{u_{d}}}=\textrm{diag}\{\sigma_{1,1},\sigma_{1,1},\dots,\sigma_{L,L}\}$. Moreover, each diagonal element follows a Circularly symmetric complex gaussian (CSCG) distribution $\mathcal{CN}(0,\frac{\rho}{L})$. Meanwhile, we model the UL and DL channels as each element of the PRV follows the CSCG distribution $\mathcal{CN}(0,\frac{g_{0}d_{d/u}^{-\alpha}}{L})$, where $g_{0}$ represents the path loss at the reference distance of 1m, $\alpha$ is the path loss exponent, and $d_{u/d}$ represents the propagation distance from BS to DL UE-$d$ or the distance from UL UE-$u$ to BS. Similarly, we can model the channel between each UL UE-$u$ and DL UE-$d$ where each PRV element follows the CSCG distribution $\mathcal{CN}(0,\frac{g_{0_{1}}d_{ud}^{-\alpha_{1}}}{L})$. Meanwhile, PRM of SI channel follows the CSCG distribution $\mathcal{CN}(0,\frac{SI}{L})$. We assume that the AoDs and AoAs are to be independent and identically distributed random variables within the interval of $[0,2\pi]$. The numerical results are averaged over 100 independent and random channel realizations. Learning rates for all neural networks are set to 0.001. Following [20], we adopt the optimal decoding order of the two-user case for UL UEs which is given as follows: $s_{1,1}<s_{2}<s_{1,2}$. Unless otherwise specified, the default simulation parameters are provided in Table I and Table II. In our simulation results, we denote our proposed ”Scenario 1” as ”FD RSMA-MA UE side”, ”Scenario 2” as ”FD RSMA-MA BS side”, and ”Scenario 3” as ”FD RSMA-MA both side”. Finally, we compare our three scenarios with the traditional FD-RSMA-FPA scheme, which represents that BS and UEs are equipped with FPAs. It should be noted that all four of these schemes consist of the same number of antennas on the BS and UE sides.
@@ V-A Simulation Settings
$N_{T}$, $N_{R}$ 4 $g_{0}$, $g_{0_{1}}$40 dB, -50 dB
$\lambda$, $A$ 0.01, $2\lambda$ $\alpha$, $\alpha_{1}$ 2.8, 3.5
@@
$\sigma_{d}^{2}$, $\sigma_{u}^{2}$, $\sigma_{U}^{2}$90 dBm $S$, $DS$ 20, $\lambda/2$
@@
V-B Convergence Analysis
Fig. 2 presents the convergence behavior of our proposed GML algorithm for the proposed FD-BS RSMA with MA. For the brevity purpose, we provide only convergence plot of ”Scenario 3” where both BS and UE are equipped with MAs. We plot the convergence of ”Scenario 3” for two power budget of BS; 30 dBm and 45 dBm respectively. It should be noted that convergence is plotted for a single channel realization only. It can be seen from the figure that when $P_{BS}=30$ dBm, our proposed algorithm converges after 300 epochs. Meanwhile, for $P_{BS}=45$ dBm, the algorithm stops increasing significantly after 500 epochs. Despite the inherent complexity and non-convexity of the formulated optimization problem, our proposed algorithm demonstrates fast convergence, typically within 300 to 500 epochs.
V-C Impact of BS Transmit Power
Fig. 3(a) illustrates the relationship between the BS transmit power and the total sum rate when the residual SI coefficient is -60 dB. It can be observed that FD RSMA-MA both side and FD-RSMA-MA BS side achieve significantly higher sum rates as the BS transmission power increases. This performance gain is attributed to the presence of MAs at the BS, which enhance adaptability in mitigating the negative effects of residual SI in FD communication. By contrast, FD-RSMA-MA UE side and FD-RSMA-FPA, which rely on FPAs at the BS, exhibit limited capability in counteracting the adverse impact of residual SI, thereby constraining their achievable sum rates. In addition to the extra-ordinary benefits achieved by deploying MA in FD-BS-RSMA system, our proposed GML scheme achieved near-optimal performance with around 98.5% accuracy. This conclusion is drawn by comparing the results of the FD-RSMA-BS-MA both side model when solving the optimization problem using the GML approach versus the Fmincon solver, which was initialized with 100 starting points to ensure solutions close to the optimal value.
The findings are further validated by Fig. 3(b) and 3(c). Specifically, Fig. 3(b) presents the performance of the DL sum rate, revealing that as the BS power budget increases, FD-RSMA-MA both side and the FD-RSMA-MA BS side experience a significant increase in the DL sum rate, while the FD-RSMA-MA UE side and FD-RSMA-FPA exhibit only a moderate increase. Similarly, Fig. 3(c) examines the impact of BS transmit power on UL sum rate. It shows that increasing the BS power results in greater SI at the BS receiver due to FD operation, which results in a lower sum rate at UL reception. Notably, FD-RSMA-MA both side and FD-RSMA-MA BS side effectively mitigate this effect through optimal MA placement, ensuring that SI has minimal impact on the UL sum rate, even as BS power increases. In contrast, FD-RSMA-MA UE side and FD-RSMA-FPA, which are equipped with FPAs at BS, suffer a sharp decline in sum rate as BS transmission power increases, further revealing the adaptability of MA in interference management. At 30 dBm of BS power budget, FD-RSMA-MA both side achieves around 9.15% performance gain over FD-RSMA-MA BS side, 73.94% gain over FD-RSMA-MA UE side and 145.32% gain over FD-RSMA-FPA in terms of total achievable sum rate. Meanwhile, at 30 dBm of BS power budget, performance gain of FD-RSMA-MA BS side over FD-RSMA-MA UE side is around 56.67% and around 120.54% gain over FD-RSMA-FPA in terms of total sum rate. Finally, FD-RSMA-MA UE side achieves around 41.03% performance gains over FD-RSMA-FPA in terms of total sum rate at 30 dBm of BS power budget.
V-D Impact of UE Transmission Power
In Fig. 4 we analyze the impact of the UE transmission power on the total sum rate, as well as the individual DL and UL sum rates, by varying the UE transmission power from 15 dBm to 23 dBm. The results indicate that as the UE transmission power increases, the total sum rate also improves. Among all evaluated schemes, FD RSMA-MA both side achieves the highest total sum rate, as it benefits from deploying MA at both the transmitter and receiver sides. This configuration fully exploits spatial diversity, leading to superior performance. In Fig. 4(a), comparing this with the scheme that considers MA only at the BS side, we observe that at lower UE power budgets, the performance gap between these two schemes remains relatively modest. However, as the UE transmission power increases, the performance gap widens. This is because FD-RSMA MA both sides, which incorporates MA at both ends, effectively mitigates SI and intra-cell interference, whereas FD-RSMA-MA BS side, which contains MA only at the BS, can address SI but struggles with intra-cell interference as the UEs relying solely on FPAs.
Furthermore, in FD-RSMA-MA UE side, where MA is implemented only at the UE side, intra-cell interference is handled more effectively compared to FD-RSMA-FPA. However, the results confirm that MA deployment on both the transmitter and receiver sides consistently outperforms all other configurations, as it maximizes interference mitigation and spatial diversity gains. At 23 dBm of UE power budget, our proposed FD-RSMA-MA on both side achieves around 15.68% performance gains over the proposed FD-RSMA-MA BS side, 64.74% performance gains over the FD-RSMA-MA UE side and 138.29% performance gains over FD-RSMA FPA in terms of total sum rate. Meanwhile, the proposed FD-RSMA-MA UE side achieves around 44.67% improvement over FD-RSMA-FPA in terms of sum rate at 23 dBm of UE transmission power budget. It should be noted that for the case of FD RSMA-FPA, when the UE transmission power is around 15 dBm to 19 dBm, it fails to produce to any feasible solutions while meeting the QoS requirements. Meanwhile, Fig. 4(b) presents only DL sum rate vs the UE transmission power. It can be seen from the figure that as we increase the transmission power of UL UEs, the DL sum rate tends to decrease a small amount, even though the BS transmission power remains the same. This is because as we increase the power budget of UEs, intra-cell interference coming from UL UE transmission towards the DL UE reception tends to increase, resulting in a lower sum rate. In this circumstance, FD-RSMA-MA both side and FD-RSMA-MA UE side seem to handle better intra-cell interference as both of them are equipped with MA in comparison to the schemes who have only FPAs at UE side (FD-RSMA-MA BS side and FD-RSMA-FPA). On the other hand, Fig. 4(c) shows the behavior of UL sum rate when we increase the UE’s transmission power. It presents that every scheme increases with a higher rate of increase exhibited by the two schemes with MAs at BS side.
@@ V-D Impact of UE Transmission Power
V-E Impact of Residual SI
@@
V-F Impact of DL UE Rate Threshold
@@ VI Conclusion
With the aim of enhancing the SE in term of sum rate of FD-BS RSMA system by handling the negative impact of residual SI and intra-cell interference, we investigat the joint optimization of transmitting and receiving beamforming vectors at BS, UL UE transmission power, common split portion of RSMA, and MA positions, which was formulated as a non-convex optimization problem. By invoking the GML technique, the non-convex optimization problem is solved in a distributed manner for the sake of the tractability of the problem. Finally, simulation results under various system parameter settings validate that deploying MA at an FD system can obtain more favorable outcomes than FPA by effectively combating residual SI and intra-cell interference. Particularly, our numerical results reveal that deploying MAs at BS is more effective than having MA only at the UE side, and the best outcome is achieved when MA is deployed at both the transmitter and receiver sides.

### v3

@@ 7    $\textbf{u}^{(0,j)}=\textbf{u}^{(0,1)}$;
10         Solve $\mathcal{P}_{4}$;
@@ 15      Apply projection if applicable as (41)-(44)
17         Solve $\mathcal{P}_{5}$;
@@ 22      Apply projection as (47)
24         Solve $\mathcal{P}_{6}$;
@@ 28    $\textbf{Z}^{*}=\textbf{Z}^{(N_{i},j)}$;
31         Solve $\mathcal{P}_{7}$;
@@ 35    $\textbf{C}^{*}=\textbf{C}^{(N_{i},j)}$;
38         Solve $\mathcal{P}_{8}$;
@@ 43      Calculate $\mathcal{L}^{j}$ as (54);
44    $\bar{\mathcal{L}}=\bar{\mathcal{L}}+\mathcal{L}^{j}$;
45    If $\mathcal{L}^{j}>MAX$
[…]
47       $\textbf{P}_{opt}=\textbf{P}^{*}$, $\textbf{W}_{opt}=\textbf{W}^{*}$, $\textbf{Z}_{opt}=\textbf{P}^{*}$$\textbf{C}_{opt}=\textbf{C}^{*}$, $\textbf{u}_{opt}=\textbf{u}^{*}$;
@@ 45    If $\mathcal{L}^{j}>MAX$
50  Calculate $\bar{\mathcal{L}}=\frac{1}{N_{o}}\sum_{j=1}^{N_{o}}\mathcal{L}^{j}$;
@@ V Simulation results
In the simulation, we assume two DL UEs and two UL UEs. The UEs are randomly and uniformly distributed in a cell centered on the FD BS located at (0,0) in an area of $200m\times 200m$. The FD BS is equipped with an equal number of transmit and receive antennas, i.e., $N_{T}=N_{R}$. Following [zhu2023modeling], we adopt the same channel where we assume that $L_{t}^{BS}=L_{r}^{BS}=L_{t}^{DL}=L_{t}^{u}=L_{t}^{d}=L_{r}^{SI}=L_{t}^{SI}\equiv L$. In this way, every PRM $\boldsymbol{\Sigma}_{d}$, $\boldsymbol{\Sigma}_{u}$, $\boldsymbol{\Sigma}_{SI}$, $\boldsymbol{\Sigma}_{u_{d}}$ becomes a diagonal matrix where $\boldsymbol{\Sigma}_{d}=\textrm{diag}\{\sigma_{1,1},\sigma_{1,1},\dots,\sigma_{L,L}\}$, $\boldsymbol{\Sigma}_{u}=\textrm{diag}\{\sigma_{1,1},\sigma_{1,1},\dots,\sigma_{L,L}\}$, $\boldsymbol{\Sigma}_{SI}=\textrm{diag}\{\sigma_{1,1},\sigma_{1,1},\dots,\sigma_{L,L}\}$, $\boldsymbol{\Sigma}_{u_{d}}=\textrm{diag}\{\sigma_{1,1},\sigma_{1,1},\dots,\sigma_{L,L}\}$. Moreover, each diagonal element follows a Circularly symmetric complex gaussian (CSCG) distribution $\mathcal{CN}(0,\frac{\rho}{L})$. Meanwhile, we model the UL and DL channels as each element of the PRV follows the CSCG distribution $\mathcal{CN}(0,\frac{g_{0}d_{d/u}^{-\alpha}}{L})$, where $g_{0}$ represents the path loss at the reference distance of 1m, $\alpha$ is the path loss exponent, and $d_{u/d}$ represents the propagation distance from BS to DL UE-$d$ or the distance from UL UE-$u$ to BS. Similarly, we can model the channel between each UL UE-$u$ and DL UE-$d$ where each PRV element follows the CSCG distribution $\mathcal{CN}(0,\frac{g_{0_{1}}d_{ud}^{-\alpha_{1}}}{L})$. Meanwhile, PRM of SI channel follows the CSCG distribution $\mathcal{CN}(0,\frac{SI}{L})$. We assume that the AoDs and AoAs are to be independent and identically distributed random variables within the interval of $0,2\pi]$. The numerical results are averaged over 100 independent and random channel realizations. Learning rates for all neural networks are set to 0.001. Following [liu2020rate], we adopt the optimal decoding order of the two-user case for UL UEs which is given as follows: $s_{1,1}<s_{2}<s_{1,2}$. Unless otherwise specified, the default simulation parameters are provided in Table I and Table II. In our simulation results, we denote our proposed ”Scenario 1” as ”FD RSMA-MA UE side”, ”Scenario 2” as ”FD RSMA-MA BS side”, and ”Scenario 3” as ”FD RSMA-MA both side”. Finally, we compare our three scenarios with the traditional FD-RSMA-FPA scheme, which represents that BS and UEs are equipped with FPAs. It should be noted that all four of these schemes consist of the same number of antennas on the BS and UE sides.
@@
$N_{T}$, $N_{R}$ 4 $g_{0}$, $g_{0_{1}}$ -40 dB, -50 dB
$\lambda$, AA 0.01, $2\lambda$ $\alpha$, $\alpha_{1}$ 2.8, 3.5
@@
$\sigma_{d}^{2}$, $\sigma_{u}^{2}$, $\sigma_{U}^{2}$ -90 dBm $S$, DSDS 20, $\lambda/2$
@@
V-A Convergence Analysis
Fig. 2 presents the convergence behavior of our proposed GML algorithm for the proposed FD-BS RSMA with MA. For brevity, we provide only the convergence plot of ”Scenario 3” where both BS and UE are equipped with MAs. We plot the convergence of ”Scenario 3” for two power budgets of BS: 30 dBm and 45 dBm, respectively. It should be noted that convergence is plotted for a single channel realization only. It can be seen from the figure that when $P_{BS}=30$ dBm, our proposed algorithm converges after 300 epochs. Meanwhile, for $P_{BS}=45$ dBm, the algorithm stops increasing significantly after 500 epochs. Despite the inherent complexity and non-convexity of the formulated optimization problem, our proposed algorithm demonstrates fast convergence, typically within 300 to 500 epochs.
V-B Impact of BS Transmit Power
Fig. 3(a) illustrates the relationship between the BS transmit power and the total sum rate when the residual SI coefficient is -60 dB. It can be observed that FD RSMA-MA on both sides and FD-RSMA-MA BS side achieve significantly higher sum rates as the BS transmission power increases. This performance gain is attributed to the presence of MAs at the BS, which enhance adaptability in mitigating the negative effects of residual SI in FD communication. By contrast, FD-RSMA-MA UE side and FD-RSMA-FPA, which rely on FPAs at the BS, exhibit limited capability in counteracting the adverse impact of residual SI, thereby constraining their achievable sum rates. In addition to the extraordinary benefits achieved by deploying MA in the FD-BS-RSMA system, our proposed GML scheme achieved near-optimal performance with around 98.5% accuracy. This conclusion is drawn by comparing the results of the FD-RSMA-BS-MA model on both sides when solving the optimization problem using the GML approach versus the Fmincon solver, which was initialized with 100 starting points to ensure solutions close to the optimal value. The findings are further validated by Fig. 3(b) and 3(c). Specifically, Fig. 3(b) presents the performance of the DL sum rate, revealing that as the BS power budget increases, FD-RSMA-MA both side and the FD-RSMA-MA BS side experience a significant increase in the DL sum rate, while the FD-RSMA-MA UE side and FD-RSMA-FPA exhibit only a moderate increase.
Similarly, Fig. 3(c) examines the impact of BS transmit power on UL sum rate. It shows that increasing the BS power results in greater SI at the BS receiver due to FD operation, which results in a lower sum rate at UL reception. Notably, FD-RSMA-MA both side and FD-RSMA-MA BS side effectively mitigate this effect through optimal MA placement, ensuring that SI has minimal impact on the UL sum rate, even as BS power increases. In contrast, FD-RSMA-MA UE side and FD-RSMA-FPA, which are equipped with FPAs at BS, suffer a sharp decline in sum rate as BS transmission power increases, further revealing the adaptability of MA in interference management. At 30 dBm of BS power budget, FD-RSMA-MA both side achieves around 9.15% performance gain over FD-RSMA-MA BS side, 73.94% gain over FD-RSMA-MA UE side and 145.32% gain over FD-RSMA-FPA in terms of total achievable sum rate. Meanwhile, at 30 dBm of BS power budget, performance gain of FD-RSMA-MA BS side over FD-RSMA-MA UE side is around 56.67% and around 120.54% gain over FD-RSMA-FPA in terms of total sum rate. Finally, FD-RSMA-MA UE side achieves around 41.03% performance gains over FD-RSMA-FPA in terms of total sum rate at 30 dBm of BS power budget.
@@ V-B Impact of BS Transmit Power
V-C Impact of UE Transmission Power
In Fig. 4 we analyze the impact of the UE transmission power on the total sum rate, as well as the individual DL and UL sum rates, by varying the UE transmission power from 15 dBm to 23 dBm. The results indicate that as the UE transmission power increases, the total sum rate also improves. Among all evaluated schemes, FD RSMA-MA both side achieves the highest total sum rate, as it benefits from deploying MA at both the transmitter and receiver sides. This configuration fully exploits spatial diversity, leading to superior performance. In Fig. 4(a), comparing this with the scheme that considers MA only at the BS side, we observe that at lower UE power budgets, the performance gap between these two schemes remains relatively modest. However, as the UE transmission power increases, the performance gap widens. This is because FD-RSMA MA on both sides, which incorporates MA at both ends, effectively mitigates SI and intra-cell interference, whereas FD-RSMA-MA BS side, which contains MA only at the BS, can address SI but struggles with intra-cell interference as the UEs rely solely on FPAs.
Furthermore, in the FD-RSMA-MA UE side, where MA is implemented only at the UE side, intra-cell interference is handled more effectively compared to FD-RSMA-FPA. However, the results confirm that MA deployment on both the transmitter and receiver sides consistently outperforms all other configurations, as it maximizes interference mitigation and spatial diversity gains. At 23 dBm of UE power budget, our proposed FD-RSMA-MA on both side achieves around 15.68% performance gains over the proposed FD-RSMA-MA BS side, 64.74% performance gains over the FD-RSMA-MA UE side and 138.29% performance gains over FD-RSMA FPA in terms of total sum rate. Meanwhile, the proposed FD-RSMA-MA UE side achieves around 44.67% improvement over FD-RSMA-FPA in terms of sum rate at 23 dBm of UE transmission power budget. It should be noted that for the case of FD RSMA-FPA, when the UE transmission power is around 15 dBm to 19 dBm, it fails to produce to any feasible solutions while meeting the QoS requirements. Meanwhile, Fig. 4(b) presents only DL sum rate vs the UE transmission power. It can be seen from the figure that as we increase the transmission power of UL UEs, the DL sum rate tends to decrease a small amount, even though the BS transmission power remains the same. This is because as we increase the power budget of UEs, intra-cell interference coming from UL UE transmission towards the DL UE reception tends to increase, resulting in a lower sum rate. In this circumstance, FD-RSMA-MA both side and FD-RSMA-MA UE side seem to handle better intra-cell interference as both of them are equipped with MA in comparison to the schemes who have only FPAs at UE side (FD-RSMA-MA BS side and FD-RSMA-FPA). On the other hand, Fig. 4(c) shows the behavior of UL sum rate when we increase the UE’s transmission power. It presents that every scheme increases, with a higher rate of increase exhibited by the two schemes with MAs on the BS side.
V-D Impact of Residual SI
@@
V-E Impact of DL UE Rate Threshold
@@ VI Conclusion
With the aim of enhancing the SE in terms of sum rate of FD-BS RSMA system by handling the negative impact of residual SI and intra-cell interference, we investigate the joint optimization of transmitting and receiving beamforming vectors at BS, UL UE transmission power, the common split portion of RSMA, and MA positions, which was formulated as a non-convex optimization problem. By invoking the GML technique, the non-convex optimization problem is solved in a distributed manner for the sake of the tractability of the problem. Finally, simulation results under various system parameter settings validate that deploying MA at an FD system can obtain more favorable outcomes than FPA by effectively combating residual SI and intra-cell interference. Particularly, our numerical results reveal that deploying MAs at BS is more effective than having MA only at the UE side, and the best outcome is achieved when MA is deployed at both the transmitter and receiver sides.